# HANS-Net: Hyperbolic Convolution and Adaptive Temporal Attention for Accurate and Generalizable Liver and Tumor Segmentation in CT Imaging

### v1

@@ Abstract
Accurate liver and tumor segmentation on abdominal CT images is critical for reliable diagnosis and treatment planning, but remains challenging due to complex anatomical structures, variability in tumor appearance, and limited annotated data. To address these issues, we introduce Hyperbolic-convolutions Adaptive-temporal-attention with Neural-representation and Synaptic-plasticity Network (HANS-Net), a novel segmentation framework that synergistically combines hyperbolic convolutions for hierarchical geometric representation, a wavelet-inspired decomposition module for multi-scale texture learning, a biologically motivated synaptic plasticity mechanism for adaptive feature enhancement, and an implicit neural representation branch to model fine-grained and continuous anatomical boundaries. Additionally, we incorporate uncertainty-aware Monte Carlo dropout to quantify prediction confidence and lightweight temporal attention to improve inter-slice consistency without sacrificing efficiency. Extensive evaluations of the LiTS dataset demonstrate that HANS-Net achieves a mean Dice score of 93.26%, an IoU of 88.09%, an average symmetric surface distance (ASSD) of 0.72 mm, and a volume overlap error (VOE) of 11.91%. Furthermore, cross-dataset validation on the 3D-IRCADb-01 dataset obtains an average Dice of 87.45%, IoU of 80.30%, ASSD of 1.525 mm, and VOE of 19.71%, indicating strong generalization across different datasets. These results confirm the effectiveness and robustness of HANS-Net in providing anatomically consistent, accurate, and confident liver and tumor segmentation.
Keywords: Liver segmentation, tumor segmentation, hyperbolic convolutions, neural representation, synaptic plasticity
@@ 1   Introduction
The liver is a vital glandular organ that performs functions of digestion, metabolism, detoxification, and immunity [1]. Liver cancer ranks among the five most frequently diagnosed cancers and constitutes the fourth leading cause of cancer-related mortality worldwide, with hepatocellular carcinoma (HCC) being the most prevalent form [2]. Precise segmentation of liver tumors is essential in the diagnosis and radiation therapy of HCC. According to the World Health Organization (WHO), cancer accounted for 8.8 million deaths in 2015, with 788,000 specifically attributed to liver cancer [3]. With increased public health awareness, the need for effective liver cancer diagnosis and treatment has intensified [4]. Historically, specialists have manually identified liver tumors, an accurate, yet tedious and labor intensive process prone to errors under substantial workload [4]. The development of computer-aided diagnostic systems has introduced intelligent, automated, or partially automated approaches to the analysis of medical images, significantly improving precision, effectiveness, and consistency in the detection and segmenting of tumors [5].
@@
Various image processing and machine learning techniques have been developed to improve liver tumor segmentation and support early cancer intervention [6]. Among these methodologies, attention-based and multi-branch models have earned significant interest, integrating sophisticated architectures such as S2DA-Net [7], EG-UNETR [8], and Generalized U-Net [9]. These models are structured to encapsulate essential semantic features in various imaging phases, increasing both the precision and the robustness of feature extraction [9]. Currently, other innovative strategies have taken advantage of multiphase fusion and multiscale representation techniques, such as PGC-Net [10] and camouflaged feature mining modules [5], which collectively advance the automated identification of complex anatomical structures, encompassing both liver and tumor regions [4].
@@
Recent developments in liver tumor segmentation have achieved significant progress, yet several challenges persist in comprehensively addressing the complexity of the task. Current liver and tumor segmentation studies frequently encounter limitations in effective management of complex and heterogeneous tumor-liver anatomies, resulting in only moderate accuracy levels. These models depend on flattened representations, which hinders their ability to accurately capture the complex spatial organization and progression of tumors across 3D slices [10]. Furthermore, they are often incapable of segmenting tumors and livers simultaneously [11]. Most methodologies overlook the enhancement of inter-slice consistency without compromising efficiency. They also do not improve adaptive features and construct continuous and fine-grained anatomical boundaries [5, 9]. Moreover, they typically lack the strategies of temporal attention necessary to capture sequential relationships in volumetric scans and face challenges in resolving fine-grained tumor boundaries due to insufficient detail encoding [4]. Finally, the absence of uncertainty estimation prevents the scope of quantifying the confidence in prediction [12].
@@
HANS-Net addresses these shortcomings through a comprehensive design. It employs hyperbolic convolutions to represent anatomical hierarchies in a non-Euclidean space, enabling accurate modeling of tumor complexity. A wavelet-inspired decomposition module captures multiscale information, while an adaptive temporal attention mechanism ensures inter-slice consistency. A synaptic plasticity module enhances feature adaptivity, and an implicit neural representation branch improves boundary smoothness. Finally, Monte Carlo dropout is used to quantify uncertainty in the predictions. HANS-Net achieves superior performance (Dice: 93.26%), surpassing state-of-the-art (SOTA) accuracy and demonstrates strong cross-dataset generalization (Dice: 87.45% on 3D-IRCADb-01), highlighting its robustness and potential for real-world clinical deployment.
@@
  • Introduces HANS-Net, the first liver tumor segmentation framework that unite hyperbolic geometry and implicit neural representations for continuous, hierarchy-aware learning in curved space.
@@
  • Integrates a bioinspired synaptic plasticity mechanism that adaptively adjusts feature weights based on activation dynamics, promoting tumor-focused learning without extra supervision.
@@
The remainder of this paper is organized as follows. Section 2 reviews the related work conducted in this area. Section 3 details our proposed HANS-Net model and its component designs. Section 4 describes the datasets, pre-processing, and experimental setup. Section 5 presents quantitative and qualitative results, including cross-dataset experiments, uncertainty estimation, and ablation studies. Section 6 discusses the implications of our findings and their relevance to current research, including potential future directions. Finally, Section 7 summarizes our findings and concludes the study.
@@ 2.1   Multi-branch and attention-based models
Automatic segmentation of liver tumors contributes significantly to supporting clinicians in making diagnosis and treatment decisions [3]. Although Convolutional neural networks (CNN) have been used effectively in medical image segmentation, they struggle to capture long-range pixel dependencies [7]. In contrast, transformer-based models require a substantial number of parameters and involve significant computational expenses [7]. Multibranch and attention-based models have been suggested as a promising alternative to overcome the drawbacks presented by both CNNs and Transformers. For example, Liu et al. [7] introduced the S2DA-Net, which incorporates a double-branched encoder with a group of multi-head cross-attention aggregation modules to improve accuracy for liver tumors. The model achieved 69.4% Dice per case and 80.0% global Dice on LiTS2017 and 3D-IRCADb, respectively. In another work, Cheng et al. [8] proposed an edge-guided segmentation network, EG-UNETR, by synthesizing TransBridge (TB) and a feature semantic enhancement module. The TB utilized multi-head position-aware attention and fusion attention to facilitate the integration of cross-level and multiscale contextual feature fusion. The model scored Dice scores of 84.45% and 85.01% on the LiTS2017 and 3D-IRCADb, respectively.
@@
Similarly, Kuang et al. [12] developed an innovative multiphase channel stacked dual attention module called MCDA, which captures essential semantic information across different phases. The model achieved a Dice score of 79.22%. On the other hand, Wang et al. [4] proposed SBCNet which incorporates a contextual encoding module to identify tumor variability through the use of a multiscale adaptive kernel, and a boundary enhancement to improve boundary perception by integrating contour learning with the Sobel operator; SBCNet scored a Dice of 82.5%. In another study, DJ et al. [9] proposed G-Unet, which is capable of incorporating other models such as CNN, residual networks (ResNets) and densely connected convolutional neural networks (DenseNet) into the general U-Net framework. G-Unet reached a global Dice of 72.9%. Finally, Hu et al. [11] introduced MSML-AttUNet, a hierarchical attention network that leverages multilayer dilated convolutions for multiscale feature extraction and expands the receptive field for richer contextual representation; it achieved a DSC of 87.74%.
@@
Attention U-Net [16] introduces spatial attention mechanisms to focus on salient features, improving boundary delineation. However, it lacks dynamic modulation based on feature activity, a gap we address using our biologically inspired Synaptic Plasticity Module. TransUNet [2] and Swin UNETR [17] leverage transformers to model global context, improving long-range dependency capture. Nevertheless, they operate in Euclidean space and fail to capture the hierarchical spatial relationships of anatomical structures; this is precisely where our Hyperbolic Convolution Layer excels by embedding features in a non-Euclidean geometric space.
@@ 2.2   Multi-Phase Fusion
Models that leverage temporal information or hierarchical representations have shown promise in advancing liver tumor segmentation. Hu et al. [18] proposed the TMPLiTS method, a trustworthy multiphase segmentation framework that fuses information from multiple temporal CT scans using a multi-expert mixture scheme, achieving a Dice score of 81.07%. Similarly, architectures such as EG-UNETR [8] and 3D U²-Net [19] utilize multiple encoding branches and nested skip connections to integrate multiscale features, allowing both local and global context modeling. Although these models are effective in aggregating spatial information, they often require extensive parameter tuning and remain limited in handling temporal coherence across 3D volumes. In contrast, HANS-Net incorporates a wavelet-inspired decomposition module to facilitate efficient and principled multiscale feature extraction. Beyond spatial scaling, it introduces an Adaptive Temporal Attention mechanism to explicitly model inter-slice dependencies in volumetric scans, an aspect frequently overlooked in slice-based or purely spatial approaches. This dual focus on multiscale structure and temporal dynamics enables HANS-Net to deliver superior segmentation accuracy while preserving anatomical consistency across frames.
@@ 2.3   Geometry and Multiscale Feature Models
Multiscale and geometry-aware architectures have gained traction for their ability to model complex anatomical variations in liver and tumor segmentation. Yang et al. [5] introduced MCFMFNet, a multiscale feature aggregation framework that utilizes multiple convolutional kernels with diverse receptive fields, combined through an element-wise fusion strategy. They also proposed a camouflaged feature mining module to focus attention on tumor regions that are often misclassified as background, achieving Dice scores of 0.8413 and 0.7365 across datasets. Similarly, You et al. [10] proposed PGC-Net, which uses a Pyramid Vision Transformer to extract multiscale regional and contour features. These features are refined via a contour-induced parallel graph reasoning mechanism that propagates complementary information through graph convolution and maps it back to the original pixel space. This method achieved Dice scores of 73.63% and 74.16% on the LiTS17 and 3DIRCADb datasets, respectively.
Beyond these transformer-based and multiscale CNN approaches, geometry-aware models such as MMF-Net [20] and ResUNet++ [21] leverage pyramid pooling, atrous spatial pyramid modules, and residual refinements to improve contextual understanding across scales. While these designs enhance segmentation performance, they are inherently limited to Euclidean space representations and lack mechanisms to model hierarchical anatomical relationships. Their reliance on stacked convolutional layers can also constrain the ability to capture long-range geometric dependencies, especially in cases of irregular tumor-liver boundaries. In contrast, HANS-Net introduces hyperbolic convolutions into the medical image segmentation domain for the first time, addressing these limitations by enabling compact and hierarchy-aware embeddings. This approach captures the latent anatomical structure more effectively, contributing to improved boundary continuity and segmentation robustness—especially important in anatomically complex regions. Combined with HANS-Net’s wavelet-inspired decomposition and temporal modeling strategies, this geometry-aware design offers a significant advancement over traditional multiscale CNNs and transformer-based methods.
@@ 3.2   Wavelet Decomposition Module
We designed the Wavelet Decomposition Module using convolutional layers to perform low-pass and high-pass filtering, followed by a 1$\times$1 convolution to fuse the resulting feature components into a combined representation [22]. The overall process can be mathematically formulated in two distinct steps.
First, the input tensor $\mathbf{x}\in\mathbb{R}^{B\times C\times H\times W}$, where $B$ is the batch size, $C$ is the number of input channels, and $H$ and $W$ are the spatial dimensions, is passed through three convolutional filters to extract the low-pass and high-pass features as shown in Equation (1):
[…]

### v2

@@ Abstract
Accurate liver and tumor segmentation on abdominal CT images is critical for reliable diagnosis and treatment planning, but remains challenging due to complex anatomical structures, variability in tumor appearance, and limited annotated data. To address these issues, we introduce Hyperbolic-convolutions Adaptive-temporal-attention with Neural-representation and Synaptic-plasticity Network (HANS-Net), a novel segmentation framework that synergistically combines hyperbolic convolutions for hierarchical geometric representation, a wavelet-inspired decomposition module for multi-scale texture learning, a biologically motivated synaptic plasticity mechanism for adaptive feature enhancement, and an implicit neural representation branch to model fine-grained and continuous anatomical boundaries. Additionally, we incorporate uncertainty-aware Monte Carlo dropout to quantify prediction confidence and lightweight temporal attention to improve inter-slice consistency without sacrificing efficiency. Extensive evaluations of the LiTS dataset demonstrate that HANS-Net achieves a mean Dice score of 93.26%, an IoU of 88.09%, an average symmetric surface distance (ASSD) of 0.72 mm, and a volume overlap error (VOE) of 11.91%. Furthermore, cross-dataset validation on the AMOS 2022 dataset obtains an average Dice of 85.09%, IoU of 76.66%, ASSD of 19.49 mm, and VOE of 23.34%, indicating strong generalization across different datasets. These results confirm the effectiveness and robustness of HANS-Net in providing anatomically consistent, accurate, and confident liver and tumor segmentation.
{IEEEkeywords}
Liver segmentation, tumor segmentation, hyperbolic convolutions, neural representation, synaptic plasticity
@@ 1 Introduction
The liver is a vital glandular organ that performs functions of digestion, metabolism, detoxification, and immunity [1]. Liver cancer ranks among the five most frequently diagnosed cancers and constitutes the fourth leading cause of cancer-related mortality worldwide, with hepatocellular carcinoma (HCC) being the most prevalent form [2]. Precise segmentation of liver tumors is essential in the diagnosis and radiation therapy of HCC. According to the World Health Organization (WHO), cancer accounted for 8.8 million deaths in 2015, with 788,000 specifically attributed to liver cancer [3, 4]. With increased public health awareness, the need for effective liver cancer diagnosis and treatment has intensified [5, 6]. Historically, specialists have manually identified liver tumors, an accurate, yet tedious and labor-intensive process prone to errors under substantial workload [5, 7]. The development of computer-aided diagnostic systems has introduced intelligent, automated, or partially automated approaches to the analysis of medical images, significantly improving precision, effectiveness, and consistency in the detection and segmentation of tumors. [8, 9, 7].
@@
Various image processing and machine learning techniques have been developed to improve liver tumor segmentation and support early cancer intervention [10, 4, 6]. Among these methodologies, attention-based and multi-branch models have earned significant interest, integrating sophisticated architectures such as S2DA-Net [11], EG-UNETR [12], Generalized U-Net [13], and ConvNeXt-2U [14]. These models are structured to encapsulate essential semantic features in various imaging phases, increasing both the precision and the robustness of feature extraction [13]. Currently, other innovative strategies have taken advantage of multiphase fusion and multiscale representation techniques, such as PGC-Net [15] and camouflaged feature mining modules [8], which collectively advance the automated identification of complex anatomical structures, encompassing both liver and tumor regions [5].
@@
Recent developments in liver tumor segmentation have made progress, yet several challenges persist in comprehensively addressing the complexity of the task. Current liver and tumor segmentation studies frequently encounter limitations in the effective management of complex and heterogeneous tumor-liver anatomies, resulting in only moderate accuracy levels. These models depend on flattened representations, which hinders their ability to accurately capture the complex spatial organization and progression of tumors across 3D slices [15]. Furthermore, they are often incapable of segmenting tumors and livers simultaneously [16]. Most methodologies overlook the enhancement of inter-slice consistency without compromising efficiency. They also do not improve adaptive features and construct continuous and fine-grained anatomical boundaries [8, 13]. Moreover, they typically lack the strategies of temporal attention necessary to capture sequential relationships in volumetric scans and face challenges in resolving fine-grained tumor boundaries due to insufficient detail encoding [5]. Finally, the absence of uncertainty estimation prevents the scope of quantifying the confidence in prediction [17].
@@
HANS-Net addresses these shortcomings through a comprehensive design. It employs hyperbolic convolutions to represent anatomical hierarchies in a non-Euclidean space, enabling accurate modeling of tumor complexity, and a wavelet-inspired decomposition module captures multiscale information. While 3D volumetric approaches naturally preserve spatial continuity, they are computationally intensive and often require extensive annotated data. In contrast, 2D slice-based methods offer computational efficiency and can effectively capture inter-slice relationships when equipped with appropriate mechanisms. Accordingly, HANS-Net integrates adaptive temporal attention (ATA) along with multi-view fusion, making the approach particularly suitable for resource-constrained clinical settings while maintaining competitive segmentation performance. A synaptic plasticity module enhances feature adaptivity, and an implicit neural representation branch improves boundary smoothness. Finally, Monte Carlo dropout is used to quantify uncertainty in the predictions. HANS-Net achieves superior performance (Dice: 93.26%), surpassing state-of-the-art (SOTA) accuracy and demonstrating strong cross-dataset generalization (Dice: 85.09% on AMOS 2022), highlighting its robustness and potential for real-world clinical deployment.
@@
  • Introduces HANS-Net, the first liver tumor segmentation framework that unites hyperbolic geometry and implicit neural representations for continuous, hierarchy-aware learning in curved space.
@@
  • Integrates a bio-inspired synaptic plasticity mechanism that adaptively adjusts feature weights based on activation dynamics, promoting tumor-focused learning without extra supervision.
@@
The remainder of this paper is organized as follows. Section 2 reviews the related work conducted in this area. Section 3 details our proposed HANS-Net model and its component designs. Section 4 describes the datasets, preprocessing, and experimental setup. Section 5 presents quantitative and qualitative results, including cross-dataset experiments, uncertainty estimation, and ablation studies. Section 6 discusses the implications of our findings and their relevance to current research, including potential future directions. Finally, Section 7 summarizes our findings and concludes the study.
@@ 2.1 Multi-branch and attention-based models
Automatic segmentation of liver tumors contributes significantly to supporting clinicians in making diagnosis and treatment decisions [3]. Although Convolutional neural networks (CNN) have been used effectively in medical image segmentation, they struggle to capture long-range pixel dependencies [11]. In contrast, transformer-based models require a substantial number of parameters and involve significant computational expenses [11]. Multibranch and attention-based models have been suggested as a promising alternative to overcome the drawbacks presented by both CNNs and Transformers, as highlighted in recent studies [11, 12, 21, 22, 17, 5, 13, 23, 16, 24, 2, 25].
@@
For example, Liu et al. [11] introduced the S2DA-Net, which incorporates a double-branched encoder with a group of multi-head cross-attention aggregation modules to improve accuracy for liver tumors. The model achieved 69.4% Dice per case and 80.0% global Dice on LiTS2017 and 3D-IRCADb, respectively. In another work, Cheng et al. [12] proposed an edge-guided segmentation network, EG-UNETR, by synthesizing TransBridge (TB) and a feature semantic enhancement module. The TB utilized multi-head position-aware attention and fusion attention to facilitate the integration of cross-level and multiscale contextual feature fusion. The model scored Dice scores of 84.45% and 85.01% on the LiTS2017 and 3D-IRCADb, respectively. Metaheuristic algorithm and multiscale hybrid attention models are also popular in other segmentation tasks [21]. For instance, Wan et al. [22] proposed MHAED-Net that builds an encoder-decoder network with the convolution normalization activation block for efficient segmentation. Similarly, Kuang et al. [17] developed an innovative multiphase channel stacked dual attention module called MCDA, which captures essential semantic information across different phases. The model achieved a Dice score of 79.22%. On the other hand, Wang et al. [5] proposed SBCNet, which incorporates a contextual encoding module to identify tumor variability through the use of a multiscale adaptive kernel, and a boundary enhancement to improve boundary perception by integrating contour learning with the Sobel operator; SBCNet scored a Dice of 82.5%. In another study, Deepak et al. [13] proposed G-Unet, which is capable of incorporating other models such as CNN, residual networks (ResNets), and densely connected convolutional neural networks (DenseNet) into the general U-Net framework. G-Unet reached a global Dice of 72.9%. In another paper, Zhang et al. [23] introduced the Decoupled Pyramid Correlation Network (DPC-Net), which uses a Pyramid Feature Encoder to extract multi-level features and decouples them along spatial and semantic dimensions. The network employed Spatial and Semantic Correlation modules to capture feature correlations, achieving a Dice score of 76.4% on the LiTS dataset. Finally, Hu et al. [16] introduced MSML-AttUNet, a hierarchical attention network that utilizes multilayer dilated convolutions for multiscale feature extraction and expands the receptive field for richer contextual representation; it achieved a Dice score of 87.74%.
@@
Oktay et al. [24] introduced Attention U-Net, a spatial attention mechanism to focus on salient features to improve boundary delineation. However, it lacks dynamic modulation based on feature activity, a gap we address using our biologically inspired Synaptic Plasticity Module. TransUNet [2] and Swin UNETR [25] utilized transformers to model global context, improving long-range dependency capture. In contrast, Chen et al. [14] proposed ConvNeXt-2U, which integrated two U-Nets with interconnected skip connections and 3D ConvNeXt blocks for segmentation. Nevertheless, these architectures in Euclidean space fail to capture the hierarchical spatial relationships of anatomical structures; this is precisely where our Hyperbolic Convolution Layer excels by embedding features in a non-Euclidean geometric space.
@@ 2.2 Multi-Phase Fusion
Recent studies [26, 12, 27] highlight the effectiveness of models that incorporate temporal information and hierarchical representations in liver tumor segmentation. To begin with, Hu et al. [26] proposed the TMPLiTS method, a trustworthy multiphase segmentation framework that fuses information from multiple temporal CT scans using a multi-expert mixture scheme, achieving a Dice score of 81.07%. Similarly, architectures such as EG-UNETR [12] and 3D U²-Net [27] utilize multiple encoding branches and nested skip connections to integrate multiscale features, allowing both local and global context modeling. Although these models are effective in aggregating spatial information, they often require extensive parameter tuning and remain limited in handling temporal coherence across 3D volumes. In contrast, HANS-Net incorporates a wavelet-inspired decomposition module to facilitate efficient and principled multiscale feature extraction. Beyond spatial scaling, it introduces an Adaptive Temporal Attention mechanism to explicitly model inter-slice dependencies in volumetric scans, an aspect frequently overlooked in slice-based or purely spatial approaches. This dual focus on multiscale structure and temporal dynamics enables HANS-Net to deliver superior segmentation accuracy while preserving anatomical consistency across frames.
@@ 2.3 Geometry and Multiscale Feature Models
Multiscale and geometry-aware architectures are increasingly explored in recent studies [8, 15, 28, 29] for their capacity to model complex anatomical variations in liver and tumor segmentation. Yang et al. [8] introduced MCFMFNet, a multiscale feature aggregation framework that utilizes multiple convolutional kernels with diverse receptive fields, combined through an element-wise fusion strategy. They also proposed a camouflaged feature mining module to focus attention on tumor regions that are often misclassified as background, achieving Dice scores of 0.8413 and 0.7365 across datasets. Similarly, You et al. [15] proposed PGC-Net, which uses a Pyramid Vision Transformer to extract multiscale regional and contour features. These features are refined via a contour-induced parallel graph reasoning mechanism that propagates complementary information through graph convolution and maps it back to the original pixel space. This method achieved Dice scores of 73.63% and 74.16% on the LiTS17 and 3DIRCADb datasets, respectively. Beyond these transformer-based and multiscale CNN approaches, geometry-aware models such as MMF-Net [28] and ResUNet++ [29] employ pyramid pooling, atrous spatial pyramid modules, and residual refinements to improve contextual understanding across scales. While these designs enhance segmentation performance, they are inherently limited to Euclidean space representations and lack mechanisms to model hierarchical anatomical relationships. Their reliance on stacked convolutional layers can also constrain the ability to capture long-range geometric dependencies, especially in cases of irregular tumor-liver boundaries. In contrast, HANS-Net introduces hyperbolic convolutions into the medical image segmentation domain for the first time, addressing these limitations by enabling compact and hierarchy-aware embeddings. This approach captures the latent anatomical structure more effectively, contributing to improved boundary continuity and segmentation robustness, especially important in anatomically complex regions. Combined with HANS-Net’s wavelet-inspired decomposition and temporal modeling strategies, this geometry-aware design offers a significant advancement over traditional multiscale CNNs and transformer-based methods.
@@ 3.2 Wavelet Decomposition Module
We designed the Wavelet Decomposition Module using convolutional layers to perform low-pass and high-pass filtering, followed by a $\times1$ convolution to fuse the resulting feature components into a combined representation [30]. The overall process can be mathematically formulated in two distinct steps. First, the input tensor $\mathbf{x}\in\mathbb{R}^{B\times C\times H\times W}$, where $B$ is the batch size, $C$ is the number of input channels, and $H$ and $W$ are the spatial dimensions, is passed through three convolutional filters to extract the low-pass and high-pass features as shown in Equation (1):
$\mathbf{F}=\text{Concat}\Big($ $~\text{Conv2D}(\mathbf{x},\mathbf{W}_{\text{low}}),$
$~\text{Conv2D}(\mathbf{x},\mathbf{W}_{\text{high}}^{h}),$
$~\text{Conv2D}(\mathbf{x},\mathbf{W}_{\text{high}}^{v})\Big)$ (1)